OpenStudy (anonymous):

Simplify & show work: 5/4x + 7/8 = 3/x

OpenStudy (anonymous):

5/4x-3/x=-7/8 5-12/4x=-7/8 -7/4x=-7/8 4x=8 x=2

OpenStudy (anonymous):

is it correct.............reply

OpenStudy (anonymous):

i am sure about the answer dont worry

OpenStudy (anonymous):

Thank you! And I was just writing it down.

OpenStudy (anonymous):

no problem welcome